# Phase space maximal entropy random walk: Langevin-like ensemble of physical trajectories

### v1

@@ Abstract
As written by statistician George Box ”All models are wrong, but some are useful”, standard diffusion derivation or Feynman path ensembles use nonphysical nowhere differentiable trajectories of infinite kinetic energy - what seems wrong, bringing question of differences if doing it more right this article is focused on. To consider ensembles of more physical trajectories, we can work in $(x,v)$ phase space like in Langevin equation with velocity controlling spatial steps, here also controlled with spatial potential $V(x)$. There will be discussed and compared 4 approaches to predict stationary probability distributions: using Boltzmann ensemble of points in space (GRW - generic random walk) or in phase space (psGRW), and analogously Boltzmann ensemble of paths in space (MERW - maximal entropy random walk) and in phase space (psMERW). They have qualitatively different predictions, hopefully allowing to decide the most appropriate for various settings by distinguishing them experimentally. Path ensembles have much stronger Anderson-like localization exactly as in quantum ground state, proposed novel in phase space has additionally increased density of velocity as in potential gradient, decreased toward barrier. While MERW is thermodynamically in agreement with quantum mechanics, psMERW suggests slight corrections from consideration of more physical trajectories, which seem natural - reduced velocities toward close barriers, increased down potential gradients.
Keywords: diffusion, phase space, Langevin equation, Schrödinger equation, maximal entropy random walk
@@ I Introduction
In standard derivation of diffusion equations, or in Feynman path integrals [1] we consider steps in space, for continuous limit using $\epsilon\propto\delta^{2}\to 0$ for $\epsilon$ being time step and $\delta$ being spatial step, what means velocity and $mv^{2}/2$ kinetic energy going to infinity ($\delta/\epsilon\to\infty$), nowhere differentiable trajectories.
@@
While such trajectories are clearly nonphysical, a basic question is if it is in an issue? It is mathematically more difficult, but doable to use ensembles of more physical trajectories - the question is if it could lead to a better agreement with experiment? The main purpose of this article is to start asking these basic but very difficult questions, by comparing 4 different approaches for a given spatial potential $V(x)$, with example predicted stationary probability distributions summarized in Fig. 1.
Figure 1: Examples of 4 types of predictions for 3 basic scenarios: 1D harmonic potential, 1D infinite well, and radial 3D $V\propto-1/r$ point electric or gravitational potential. Top: predicted $\rho(x)$ spatial stationary probability densities - for GRW and psGRW it is Boltzmann distribution in space. For MERW and QM it is Boltzmann path ensemble in space - $\rho\propto\psi^{2}$ for satisfying (Schrödinger) eigenequation $E\psi(x)=(-\Delta_{x}+V(x))$ for minimal energy $E$. For proposed psMERW it is Boltzmann path ensemble in phase space - as for MERW but for $E\psi(x,v)=(-\Delta_{v}-2v\cdot\nabla_{x}+V(x)+\|v\|^{2})\,\psi(x,v)$ eigenequation (found with Wolfram Mathematica ”NDEigensystem”). Center: stationary probability density in phase space for psGRW (upper) and psMERW (lower). Below are its chosen cross-sections for positions shown with colors. For infinite potential well, near barriers only opposite velocities should be allowed - what is seen here only for psMERW. For harmonic potential we see increased density of velocity down the potential gradient.
@@
To consider ensembles of more physical trajectories, as in Langevin equation [2] we can go to $(x,v)$ phase space - include finite velocity $v$ in evolved state, which evolves randomly and controls evolution in space $(x_{t+1}-x_{t}\propto v_{t})$. To get various behaviors which could be compared experimentally, there is included general spatial potential $V(x)$.
@@
Another basic crucial question is if we should base on maximizing local entropy production GRW (generic random walk) with walker considering single steps, what leads to Boltzmann ensemble in space, or maybe better maximizing mean entropy production MERW (maximal entropy random walk) using Boltzmann ensemble of entire paths - like random walk along Ising sequence, or ”Wick-rotated” Feynman path ensembles, leading to the same stationary probability distribution as quantum ground state, with much stronger Anderson-like localization property [3].
@@
Neutrons are another objects confirmed to have quantum ground state stationary probability distribution [6] as predicted by both MERW and QM. It brings a difficult general question which approach should be used in which case, e.g. in diffusion of molecules, solitons, or dust halos in astronomical settings.
Figure 2: Three basic philosophies to predict stationary probability distributions. Left: in stochastic picture we guess stochastic propagator, usually maximizing entropy separately for each step (GRW), what leads to Boltzmann ensemble of positions in space. Right: in ergodic picture we assume we fully control deterministic evolution. Center: in contrast, in thermodynamics, statistical physics we are not so certain we know such details. Instead, we assume the safest statistical model accordingly to Jaynes principle: maximizing entropy. For each point we consider Boltzmann ensemble of possible paths from it, their statics give MERW stochastic propagator. It is nonlocal (depends on the entire situation), but effective model - used only by us to predict the most likely distributions, not used by the object itself - governed by some complex unknown evolution. In this article we expand MERW philosophy to phase space, using ensemble of more physical trajectories.
@@
Beside Langevin equation, related approaches are e.g. phase-space formulation of quantum mechanics [7], or Feynman path ensemble in phase space [8]. However, they don’t actually use velocities to choose spatial step, leading to different equations.
While for psGRW the stationary densities as just products of Boltzmann distributions for space and velocity, for psMERW they interact - leading to looking obvious physical properties as we can see in Fig. 1: in gradient of potential there is increased density of velocity down this gradient, near potential barriers there is decreased density of velocity toward it.
Unfortunately the formulas become more complicated, requiring to solve functional eigenequations, for which analytical solutions could be found rather only for very simple cases. Fortunately there are available tools to do it numerically, like the used NDEigensystem function of Wolfram Mathematica.
Section II introduces to GRW and MERW philosophies, Section III contains their proposed extension to phase space. This is initial version of article, hopefully stimulating further research in this direction, especially toward experimental distinguishing of the discussed 4 approaches.
Figure 3: Gathered GRW, MERW formulas for graph defined by matrix $M$.
@@ II Generic and maximal entropy random walk
This section briefly introduces to GRW/MERW [^1] MERW introduction: https://community.wolfram.com/groups/-/m/t/2924355 random walks summarized in Fig. 2, 3, for details see e.g. [3, 9].
While further we will use lattice for continuous limit, let us start with general random walk on a graph given by adjacency matrix $M$: $M_{ij}=0$ if there is no edge between $i$ and $j$, otherwise it is 1 or generally has some weight, for Boltzmann ensemble useful to parameterize $M_{ij}=\exp(-\beta E_{ij})$ for $E_{ij}$ being step energy and thermodynamical $\beta=1/k_{B}T$.
@@
and using only allowed transitions: $M_{ij}=0\Rightarrow S_{ij}=0$, what is equivalent with assigning $E_{ij}=\infty$ energy to this transition.
@@ II-A Maximal entropy random walk (MERW)
The basic tool to calculate its stationary distribution and propagator is noticing that $M^{l}$ matrix power for $M_{ij}=\exp(-\beta E_{ij})$, combinatorially contains sum of Boltzmann path ensemble:
$(M^{l})_{ij}=\sum_{\gamma:\gamma_{0}=i,\gamma_{l}=j}\exp(-\beta E(\gamma))\qquad\textrm{for}\ \ E(\gamma)=\sum_{i=1}^{l}E_{\gamma_{i-1}\gamma_{i}}$ (2)
We are interested in $l\to\infty$ limit, for which using natural assumption of connected and acyclic graph, the Frobenius-Perron theorem says there is a single dominant eigenvalue $\lambda$, allowing to use the dominant eigenvector $\psi$ in the limit, amplitude as result of propagator from past or future infinity:
$M^{l}\propto\lambda^{l}\psi\psi^{T}\equiv\lambda^{l}|\psi\rangle\langle\psi|\qquad\textrm{for}\quad\lambda\to\infty$ (3)
@@
$\rho^{MERW}_{i}\propto\lim_{\l\to\infty}(M^{l})_{ji}(M^{l})_{ik}\propto(\psi_{i})^{2}$ (5)
Figure 4: Diagrams for derivation of MERW formulas from path ensembles, exploiting that $(M^{l})_{ij}$ combinatorially contains summation over uniform/Boltzmann ensemble of length $l$ paths from $i$ to $j$. To get stochastic propagator $S^{MERW}_{ij}$, we can consider length $l$ paths from vertex $i$, and look at distribution of their first step for $l\to\infty$ limit. To get stationary probability distribution we can multiply such two propagators: $\rho^{MERW}_{i}\propto\lim_{l\to\infty}(M^{l})_{ji}(M^{l})_{ik}$.
II-B Lattice and continuous limit to Schrödinger equiation
Let us discretize space $\mathbb{R}^{D}\ni x=\delta\tilde{x}$ for $\tilde{x}\in\mathbb{Z}^{D}$ and time $\mathbb{R}\ni t=\epsilon\tilde{t}$ for $\tilde{t}\in\mathbb{Z}$. For continuous limit $\epsilon,\delta\to 0$. Also disretize spatial potential: $V_{\tilde{x}}\equiv V(\delta\tilde{x})$. For simplicity let us work in $D=1$, with mentioned generalization for $D>1$.
@@ II-B1 GRW leading to Boltzmann space ensemble
$M_{\tilde{x},\tilde{x}+s}=\exp(-\beta((V_{\tilde{x}}+V_{\tilde{x}+s})/2)\quad\textrm{for }s\in\{-1,0,1\}$ (6)
as it symmetric, GRW stationary probability distribution is $\propto M_{\tilde{x},\tilde{x}-1}+M_{\tilde{x},\tilde{x}}+M_{\tilde{x},\tilde{x}+1}$. Assuming continuous potential $V(x)$, in continuous limit we get just Boltzmann ensemble in space:
@@
II-B2 MERW leading to Schrödinger equiation
@@
Due to change of sign, eigenvalue maximization becomes minimization of $(3-\lambda)/(3\beta\epsilon)$, which in continuous limit should converge to some energy $E$. Assuming $\epsilon=\delta^{2}$ the difference term will lead to laplacian, hence in continuous limit we can write the eigenequation as stationary Schrödinger equitation searching for the lowest energy ground state:
$E\psi=-C\Delta\psi+V\psi\qquad\rho^{MERW}(x)\propto(\psi(x))^{2}$ (9)
@@ III Random walks in phase space
$m\frac{dx}{dt}=v\qquad m\frac{dv}{dt}=-\Gamma v+\eta$
@@
Like discussed previously, performing continuous limit of $M_{(\tilde{x},\tilde{v}),(\tilde{x}+\tilde{v},\tilde{v}\pm 1)}\sim\exp(-\bar{V}(x,v))$ for phase space GRW (psGRW), the stationary probability distribution would be just:
$\rho^{psGRW}(x,v)\propto\exp(-\beta(V(x)+mv^{2}/2))$ (11)
III-A MERW continuous limit in phase space
Let us now discretize space $\mathbb{R}^{D}\ni x=\delta\tilde{x}$ for $\tilde{x}\in\mathbb{Z}^{D}$ and time $\mathbb{R}\ni t=\epsilon\tilde{t}$ for $\tilde{t}\in\mathbb{Z}$. For continuous limit we will take $\epsilon,\delta\to 0$. Discretized $\tilde{v}\in\mathbb{Z}^{D}$ velocity for $\tilde{x}\to\tilde{x}+\tilde{v}$ transition in single step corresponds to $v=\frac{\delta}{\epsilon}\tilde{v}\in\mathbb{R}^{D}$ real velocity. As previously we calculate for $D=1$, then mention for general $D$.

### v4

# Phase space maximal entropy random walk: Langevin-like ensembles of physical trajectories

@@ Abstract
As written by statistician George Box ”All models are wrong, but some are useful”, standard diffusion derivation or Feynman path ensembles use nonphysical infinite velocity/kinetic energy nowhere differentiable trajectories - what seems wrong, might be only our approximation to simplify mathematics. This article introduces some basic tools to investigate this issue. To consider ensembles of more physical finite velocity trajectories, we can work in $(x,v)$ phase space like in Langevin equation with velocity controlling spatial steps, here also controlled with spatial potential $V(x)$. There are derived and compared 4 approaches to predict stationary probability distributions: using Boltzmann ensemble of steps/points in space (GRW - generic random walk) or in phase space (psGRW), and analogously Boltzmann ensemble of paths in space (MERW - maximal entropy random walk) and in phase space (psMERW), also generalized to Lévy flights. Path ensembles generally have much stronger Anderson-like localization, MERW has stationary distribution exactly as quantum ground state. Proposed novel MERW in phase space has some slight differences, which might be distinguished experimentally. For example for 1D infinite potential well: $\rho=1$ stationary distribution for step ensemble, $\rho\sim\sin^{2}$ for path ensemble (as in QM), and $\rho\sim\sin$ for proposed smooth path ensembles - more frequently approaching the barriers due to randomly gained velocity.
Keywords: diffusion, phase space, Langevin, Schrödinger equation, tunneling, maximal entropy random walk, Lévy flights
@@ I Introduction
In standard derivation of diffusion equation, or in Feynman path integrals [1], we consider steps in space, in continuous limit using $\epsilon\propto\delta^{2}\to 0$ for $\epsilon$ being temporal step and $\delta$ being spatial step - because width of Gaussian grows with square root of the number of steps. It means velocity and $mv^{2}/2$ kinetic energy goes to infinity ($v=\delta/\epsilon\to\infty$), nowhere differentiable trajectories.
@@
While such infinite velocity trajectories seem clearly nonphysical, a basic question is if physics really use them? Maybe only we use them to simplify mathematics? It is mathematically more difficult, but doable to use ensembles of more physical trajectories - bringing question if they could lead to a better agreement with experiment? The main purpose of this article is to start asking these basic but very difficult questions, by comparing 4 different approaches for a given spatial potential $V(x)$. Fig. 1 summarizes examples of predicted stationary probability distributions. For example uniform vs $\sim\sin^{2}$ vs $\sim\sin$-like behavior near a barrier from the infinite well case might be distinguishable. A dynamical example as tunneling with concrete final velocities is shown in Fig. 2, and further in Section V - as expected getting exponential probability decrease with barrier width, but also dependence from the initial velocity, intuitively required but neglected in standard approaches. Basic code is available in the link on the right.
Figure 1: While (Jaynes) maximal entropy principle says we should use Boltzmann ensembles, being at heart of statistical physics, there are various ways to apply it, e.g. steps, paths and smooth paths ensembles. They lead to different predictions, this way bringing difficult question: which one is the most appropriate for various scenarios. The novel smooth path ensembles are obtained by going to phase space - with velocity controlling spatial step, and being randomly modified: $(\tilde{x},\tilde{v})\to(\tilde{x}+\tilde{v},\tilde{v}\pm 1)$. Both MERW and psMERW have stationary distributions being ground states of Schrödinger equation, the latter using its written phase space version: with Laplacian $\Delta_{v}$ for velocity coming from $\tilde{v}\to\tilde{v}\pm 1$ transition. There are shown example trajectories for 1D infinite potential well, generated with Mathematica code in https://community.wolfram.com/groups/-/m/t/3124320. We can see smooth paths are more likely to reach barriers thanks to randomly gained velocity $v$, which has distribution close to Gaussian. In the bottom part there are shown stationary probability distributions: for all approaches in space (up) and for psMERW in phase space (down), for 3 popular potentials: 1D harmonic potential, 1D infinite well as above, and radial 3D $V\propto-1/r$ point electric or gravitational potential.
@@
To consider ensembles of more physical trajectories, as in Langevin equation [2] we can go to $(x,v)$ phase space - include finite velocity $v$ in evolved state, which changes randomly and controls evolution in space $(x_{t+1}-x_{t}\propto v_{t})$. To get various behaviors which could be compared experimentally, there is included general spatial potential $V(x)$.
@@
Another basic crucial question is if we should base on maximizing local entropy production GRW (generic random walk) with walker considering single steps, what leads to Boltzmann ensemble in space. Or maybe better use maximizing mean entropy production MERW (maximal entropy random walk) using Boltzmann ensemble of entire paths - like random walk along Ising sequence, or ”Wick-rotated” Feynman path ensembles, Euclidean quantum mechanics [3] - mathematically leading to the same stationary probability distribution as quantum ground state, with much stronger Anderson-like localization property [4].
@@
Neutrons are another objects confirmed to have quantum ground state stationary probability distribution [7] as predicted by both MERW and QM. Also for ”walking droplets”: classical objects with wave-particle duality, there were experimentally observed QM-like statistics [8]. They brings a difficult general question which approach should be used in which case, e.g. for diffusion of molecules, solitons, or dust halos in astronomical settings.
@@
Beside Langevin equation, related approaches are e.g. phase-space formulation of quantum mechanics [9], or Feynman path ensemble in phase space [10]. However, they do not actually use velocities to choose spatial step, leading to different equations. While here we consider Boltzmann path ensembles in phase space, for Feynman it can be found in [11] discussed further.
Figure 2: Tunneling comparison for dynamic scenario: example of discrete numerical evolution of probability distribution for MERW and psMERW on size 200 lattice in space (21 in velocity) and infinite potential well with additional symmetric potential barrier in the center. The parameters were chosen to get similar behavior - still obtaining visible qualitative differences, like boundary behavior (randomly gained velocity in psMERW makes it easier to reach barriers), differences in warmup, and widths of distributions. We can see velocity asymmetry during warmup - toward barrier have faster entropy growth. While MERW/QM tunneling velocity is a controversial question by some claimed to use superluminal ([12, 13]) - we see it rather diffuses through barrier, in phase space we get concrete finite velocities (like electrons in [14]) - randomly gained velocity to pierce the barrier. Fig. 7 shows simulations of transition probability dependence from both barrier width, but also initial velocity.
Unfortunately formulas for such more physical trajectories become more complicated, requiring to solve functional eigenequations, for which analytical solutions could be found rather only for very simple cases. Fortunately there are available tools to do it numerically, like the used NDEigensystem function of Wolfram Mathematica.
Section II introduces to GRW and MERW philosophies including continuous limit. Section III contains their proposed extension to phase space. This is early version of article, hopefully stimulating further research in this direction, especially toward experimental distinguishing of the discussed 4 approaches. Section IV briefly discusses expansion to infinite variance steps: ensembles of Lévy flights in space and phase space. Section V discusses tunneling in this approach, especially dependence from initial velocity.
@@ II Generic and maximal entropy random walk
This section briefly introduces to GRW/MERW [^1] MERW introduction: https://community.wolfram.com/groups/-/m/t/2924355 random walks summarized in Fig. 3, 4, for details see e.g. [4, 15].
While further we will use lattice for continuous limit, let us start with general random walk on a graph given by adjacency matrix $M$: $M_{ij}=0$ if there is no edge between $i$ and $j$, otherwise it is 1, or for generality it has some weight here. For Boltzmann path ensemble it is natural to parameterize:
$M_{ij}=\exp(-\beta V_{ij})\qquad(\textrm{e.g.}\ V_{ij}=(V_{i}+V_{j})/2)$ (1)
for $V_{ij}$ being energy of $i\to j$ step, and $\beta=1/k_{B}T$ in thermodynamics for $T$ temperature, for quantum-like statistics can be related with Planck’s $\hbar$.
@@
and using only allowed transitions: $M_{ij}=0\Rightarrow S_{ij}=0$, what is equivalent with assigning $V_{ij}=\infty$ energy to this transition.
@@ II-A Maximal entropy random walk (MERW)
Figure 3: Three basic philosophies to predict stationary probability distributions. Left: in stochastic picture we basically guess stochastic propagator, usually maximizing entropy separately for each step (GRW), what leads to Boltzmann ensemble of positions in space. Right: in ergodic picture we assume we fully control deterministic evolution. Center: in contrast, in thermodynamics, statistical physics we do not assume knowing such details. Instead, we assume the safest statistical model accordingly to Jaynes’ principle: maximizing entropy. It leads to: for each point consider Boltzmann ensemble of possible paths from it, their statics give MERW stochastic propagator. It is nonlocal (depends on the entire situation), but effective model - used only by us to predict the most likely distributions, not used by the object itself - governed by some complex unknown evolution. In this article we expand MERW philosophy to phase space, using Boltzmann ensemble of more physical trajectories.
Figure 4: Gathered GRW, MERW formulas for graph defined by matrix $M$.
The basic tool to calculate its stationary distribution and propagator is noticing that for $M_{ij}=\exp(-\beta V_{ij})$, matrix power $M^{l}$ combinatorially contains sum of Boltzmann path ensemble:
$(M^{l})_{ij}=\sum_{\gamma:\gamma_{0}=i,\gamma_{l}=j}\exp(-\beta E(\gamma))\qquad\textrm{for}\ \ E(\gamma)=\sum_{i=1}^{l}V_{\gamma_{i-1}\gamma_{i}}$ (3)
We are interested in $l\to\infty$ limit, for which using natural assumption of connected and acyclic graph, the Frobenius-Perron theorem says there is a single dominant eigenvalue $\lambda$, allowing to use $\phi,\psi$ dominant eigenvectors in the limit:
$M^{l}\propto\lambda^{l}\psi\phi^{T}\equiv\lambda^{l}|\psi\rangle\langle\phi|\qquad\textrm{for}\quad l\to\infty$ (4)
$M\psi=\lambda\psi\qquad\phi^{T}M=\lambda\phi^{T}\quad\textrm{maximizing}\ |\lambda|$
@@
$\rho^{MERW}_{i}\propto\lim_{\l\to\infty}(M^{l})_{ji}(M^{l})_{ik}\propto\phi_{i}\psi_{i}$ (6)
This way $\phi$ is distribution at the end of past half-paths, $\psi$ at the end of future half-paths. They come from $M^{l}$ or $(M^{T})^{l}$ for $l\to\infty$ propagators from minus/plus infinity, differing by time direction. To get some position we need to randomly get it from both direction, so its probability is $\rho_{i}\propto\phi_{i}\psi_{i}$ product of two probabilities. For (time) symmetric $M=M^{T}$ both eigenvectors are equal: $\phi=\psi$, getting $\rho_{i}\propto\psi_{i}^{2}$ Born rule.
Figure 5: Diagrams for derivation of MERW formulas from path ensembles, exploiting that for $M_{ij}=\exp(-\beta V_{ij})$, matrix power $(M^{l})_{ij}\approx\lambda^{l}\psi\phi^{T}$ combinatorially contains summation over uniform/Boltzmann ensemble of length $l$ paths from $i$ to $j$. To get stochastic propagator $S^{MERW}_{ij}$, we can consider length $l$ paths from vertex $i$, and look at distribution of their first step for $l\to\infty$ limit. To get stationary probability distribution we can multiply such two propagators: $\rho_{i}\propto\lim_{l\to\infty}(M^{l})_{ji}(M^{l})_{ik}\propto\phi_{i}\psi_{i}$.
Figure 6: Boltzmann path ensembles: $\psi$ is probability distribution at the end of future half-paths, $\phi$ at the end of past half-paths. They can be viewed as results of propagator from plus/minus infinity: $M^{l}$, $(M^{T})^{l}$ for $l\to\infty$. To get stationary probability distribution in the center of full paths we need to randomly get it from both directions: $\rho=\phi\psi$ Born rule. While for (time) symmetric $M=M^{T}$ both are equal $\phi=\psi$, in the discussed phase space case, time symmetry inverses velocity - on the right there are 3 cases as in Fig. 1, now with shown $\psi(x,v)=\phi(x,-v)$, which differ by inversion of velocity, making $\rho=\phi\psi$ symmetric. This symmetry is required as inverting all paths in their Feynman ensemble (including velocities), we get the same ensemble. While here for Boltzmann path ensembles we need two real eigenfunctions, in Feynman path ensembles they are combined in real/imaginary parts of single complex eigenfunction [11].
II-B Lattice and MERW continuous limit to Schrödinger equation
Let us discretize space $\mathbb{R}^{D}\ni x=\delta\tilde{x}$ for $\tilde{x}\in\mathbb{Z}^{D}$ and time $\mathbb{R}\ni t=\epsilon\tilde{t}$ for $\tilde{t}\in\mathbb{Z}$. For continuous limit $\epsilon,\delta\to 0$. Also discretize spatial potential: $V_{\tilde{x}}\equiv V(\delta\tilde{x})$. For simplicity let us work in $D=1$, with mentioned generalization for $D>1$.
@@ II-B1 GRW leading to Boltzmann space ensemble
$M_{\tilde{x},\tilde{x}+s}=\exp(-\beta(V_{\tilde{x}}+V_{\tilde{x}+s})/2)\quad\textrm{for }s\in\{-1,0,1\}$ (7)
as it is symmetric, GRW stationary probability distribution is $\rho_{\tilde{x}}\propto M_{\tilde{x},\tilde{x}-1}+M_{\tilde{x},\tilde{x}}+M_{\tilde{x},\tilde{x}+1}$. Assuming continuous potential $V(x)$, in the limit we get just Boltzmann ensemble in space:
@@
II-B2 MERW leading to Schrödinger ground state
@@
Due to change of sign, maximization of eigenvalue $|\lambda|$ becomes minimization of $(3-\lambda)/(3\beta\epsilon)$, which in continuous limit should converge to some energy $E$. Assuming $\epsilon=\delta^{2}$, the difference term will lead to Laplacian, hence in continuous limit we can write the eigenequation as stationary Schrödinger equitation searching for the lowest energy $E$ ground state:
$E\psi=-C\Delta\psi+V\psi\qquad\quad\rho^{MERW}(x)\propto(\psi(x))^{2}$ (10)
@@ III Random walks in phase space
$m\frac{dx}{dt}=v\qquad\qquad m\frac{dv}{dt}=-\Gamma v+\eta$
@@
III-A GRW phase space continuous limit (psGRW)
Discretizing $\mathbb{Z}\ni\tilde{x}=x/\delta$ position, and $\mathbb{Z}\ni\tilde{v}=v/\zeta$ velocity, GRW as Boltzmann ensemble of single steps e.g. $s,s^{\prime}\in\{-1,0,1\}$ can be chosen as:
$S_{(\tilde{x},\tilde{v}),(\tilde{x}+\tilde{v},\tilde{v}+s)}=\frac{\exp(-\beta(V(\delta\tilde{x})+\frac{1}{2}m(\zeta(\tilde{v}+s))^{2}))}{\sum_{s^{\prime}}\exp(-\beta(V(\delta\tilde{x})+\frac{1}{2}m(\zeta(\tilde{v}+s^{\prime}))^{2}))}$
The spatial potential contributions $V(\delta\tilde{x})$ cancel out - leaving random walk of velocity alone, in $mv^{2}/2$ kinetic energy acting as harmonic potential for velocity.
Therefore, while $M$ is not symmetric due to velocity inversion in $M^{T}$, for velocity alone it can be made symmetric, leading to $\rho(v)\propto\exp(-\beta mv^{2}/2)$ Gaussian stationary probability distribution for velocity, independent from position.
For such position independent symmetric Gaussian velocity distributions, in infinitesimal limit we can get symmetric $M$ among positions, allowing to conclude psGRW stationary probability distributions as Boltzmann distribution in the phase space:
$\rho^{psGRW}(x,v)\propto\exp(-\beta(V(x)+mv^{2}/2))\rho$ (12)
III-B MERW phase space continuous limit (psMERW)
III-B1 Derivation
Let us discretize space $\mathbb{R}^{D}\ni x=\delta\tilde{x}$ for $\tilde{x}\in\mathbb{Z}^{D}$ and time $\mathbb{R}\ni t=\epsilon\tilde{t}$ for $\tilde{t}\in\mathbb{Z}$. For continuous limit we will take $\epsilon,\delta\to 0$. Discretized $\tilde{v}\in\mathbb{Z}^{D}$ velocity for $\tilde{x}\to\tilde{x}+\tilde{v}$ transition in single step corresponds to $v=\frac{\delta}{\epsilon}\tilde{v}\in\mathbb{R}^{D}$ real velocity. As previously we calculate for $D=1$, then mention for general $D$.